OpenStudy (sbentleyaz):

Given f '(x) = (1 - x)(4 - x), determine the intervals on which f(x) is increasing or decreasing.

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle f(x)}\) is decreasing iff \(\color{black}{\displaystyle f'(x)<0}\).

OpenStudy (solomonzelman):

Explanation to the above: As you know \(\color{black}{\displaystyle f'(x)}\) is the slope of \(\color{black}{\displaystyle f(x)}\), and you also know that the function is decreasing if this slope (or \(\color{black}{\displaystyle f'(x)}\)) is negative. This is why we set \(\color{black}{\displaystyle f'(x)<0}\).

OpenStudy (solomonzelman):

With similar logic, if you want to start off from finding the intervals where \(\color{black}{\displaystyle f(x)}\) is increasing, then you set \(\color{black}{\displaystyle f'(x)>0}\).

OpenStudy (solomonzelman):

So, for example if I had: \(\color{black}{\displaystyle f'(x)=(x+3)(x-2)}\) Then, I set \(\color{black}{\displaystyle f'(x)>0}\) \(\color{black}{\displaystyle (x+3)(x-2)>0}\) The above is true if both parenthesis are positive or both negative. Therefore, \(\color{black}{\displaystyle x\in(-\infty,\,-3)}\) and \(\color{black}{\displaystyle x\in(2,\,\infty)}\) are the intervals where \(\color{black}{\displaystyle f}\) is increasing. You know that at \(\color{black}{\displaystyle x=2,\,-3}\) the slope is 0. So, therefore, you have the following interval where \(\color{black}{\displaystyle f}\) decreases: \(\color{black}{\displaystyle x\in(-3,\,2)}\) (By excluding other values of the function.)

OpenStudy (solomonzelman):

So, the take-away here is: \(\color{blue}{\small \bf \displaystyle [1] }\) \(\color{black}{\displaystyle f(x) }\) is increasing on \(\color{black}{\rm \displaystyle I }\) IFF \(\color{black}{\displaystyle f'(x)>0 }\) on \(\color{black}{\rm \displaystyle I }\). \(\color{blue}{\small \bf \displaystyle [2] }\) \(\color{black}{\displaystyle f(x) }\) is decreasing on \(\color{black}{\rm \displaystyle I }\) IFF \(\color{black}{\displaystyle f'(x)<0 }\) on \(\color{black}{\rm \displaystyle I }\).